OpenStudy (anonymous):

How do I solve ln x + ln x=0? Was never taught natural log or log definitions ever sorry;( Please help!!

OpenStudy (anonymous):

ln x + ln x = 0 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2 ln(x) = 0

OpenStudy (phi):

you can combine ln(x) + ln(x) using ln(a*b)= ln(a)+ln(b) as a first step

OpenStudy (phi):

then make each side the exponent to the base e

OpenStudy (anonymous):

Okay but once I make each side the base of e what does e represent? Or how do I solve for e because I was never taught that either:( sorry probably sound really stupid to you guys right now.

OpenStudy (phi):

you need to know ln (natural log) is the same thing as log base e (e=2.718.....) there is this rule \[ e^{\ln(a)} = a\] for your problem ln(x)+ln(x) = 0 ln(x*x) = 0 (using the rule posted above) ln(x^2) = 0 (another way to write x*x) now make each side the exponent of base e \[ e^{\ln(x^2)} = e^0 \] using the rule \[ e^{\ln(a)} = a\] this is the same as \[ x^2 = e^0\] any number to the zero power is 1 (important fact) \[ x^2= 1 \\ x= \sqrt{1}= 1\] you only use the positive root, because we do not take logs of negative numbers

OpenStudy (phi):

see http://www.khanacademy.org/math/algebra/logarithms-tutorial/logarithm_basics/v/logarithms for video(s) on logs... they might help.

OpenStudy (anonymous):

Thank you SO much!! I really appreciate it

OpenStudy (phi):

Here is Khan's whole list of videos on logs http://www.khanacademy.org/math/algebra/logarithms-tutorial It would take a big bag of popcorn to get through all of them, but watching a few might be good

OpenStudy (phi):

Here is wolfram's solution http://www.wolframalpha.com/input/?i=ln%28x%29%2Bln%28x%29%3D+0

OpenStudy (anonymous):

thank you so much I really appreciate it because I had no idea